### v1

# Holistic Explainable AI (H-XAI): Extending Transparency Beyond Developers in AI-Driven Decision Making

Abstract
Current eXplainable AI (XAI) methods largely serve developers, often focusing on justifying model outputs rather than supporting diverse stakeholder needs. A recent shift toward Evaluative AI reframes explanation as a tool for hypothesis testing, but still focuses primarily on operational organizations. We introduce Holistic-XAI (H-XAI), a unified framework that integrates causal rating methods with traditional XAI methods to support explanation as an interactive, multi-method process. H-XAI allows stakeholders to ask a series of questions, test hypotheses, and compare model behavior against automatically constructed random and biased baselines. It combines instance-level and global explanations, adapting to each stakeholder’s goals, whether understanding individual decisions, assessing group-level bias, or evaluating robustness under perturbations. We demonstrate the generality of our approach through two case studies spanning six scenarios: binary credit risk classification and financial time-series forecasting. H-XAI fills critical gaps left by existing XAI methods by combining causal ratings and post-hoc explanations to answer stakeholder-specific questions at both the individual decision level and the overall model level.
1 Introduction
Explainable AI (XAI) has long promised to make AI models more transparent and trustworthy. But despite the variety of methods available, most XAI tools are still designed for developers or model builders, not for the full range of stakeholders who interact with or are affected by AI systems (Mueller et al. 2019; Bhatt et al. 2020; Deshpande and Sharp 2022). Figure 1(b) illustrates three such key stakeholder groups: individual users (e.g., customers), operational organizations (e.g., data scientists), and regulatory bodies (e.g., policy-makers).
Consider the task of credit risk assessment. An applicant might reasonably ask, “Would my loan have been approved if I had requested a smaller amount?” A data scientist may wonder, “Which applicant features drive approval decisions on average?” A regulator might want to assess whether a model systematically favors certain demographic groups. These questions are all valid, but they differ in scope, require different types of explanations, and cannot be answered by a single method. Yet, current XAI methods are still geared toward justifying predictions rather than helping stakeholders explore models interactively. They typically do not support toggling between hypotheses (Hoffman et al. 2023b), testing counterfactuals (Miller 2023), or surfacing model limitations (Suresh et al. 2021). Most generate one-shot outputs, assuming explanation is a destination, not a process. But as any user who’s puzzled over a model’s behavior can attest, explanation is rarely a one-off affair (Hoffman et al. 2023a). Past work on black-box model rating has taken some steps toward user-centered analysis. (Srivastava and Rossi 2018) proposed a two-step method to rate AI services for bias, along with a tool for exploring trade-offs between accuracy and gender fairness in translation systems (Bernagozzi et al. 2021b). This approach, designed for settings where training data is unavailable, has since been applied to more complex systems like chatbots (Srivastava et al. 2020) and search engines (Tian et al. 2023). These efforts help communicate how models behave in the wild, but they still stop short of providing explanations aligned with stakeholder questions.
Causal fairness methods go further. They attempt to isolate the influence of protected attributes on the model’s predictions (Carey and Wu 2022). This aligns well with human reasoning: people often want to know whether changing a protected attribute, while holding everything else constant, would change the model’s decision. Building on this foundation, (Lakkaraju 2022; Srivastava et al. 2024) introduced a causal rating framework to evaluate black-box AI systems across domains, from sentiment analysis (Lakkaraju, Srivastava, and Valtorta 2024) to time-series forecasting (Lakkaraju et al. 2024, 2025). This approach quantifies both statistical and confounding bias and examines how input perturbations affect outputs. But while it helps isolate causal effects, it falls short when it comes to answering instance-level questions that matter in practice.
In this paper, we propose a unified framework, Holistic-XAI (H-XAI), that integrates traditional XAI techniques with rating-driven explanations (RDE). RDE is well-suited for questions about model robustness, bias, and the effects of input perturbations, especially when those questions require comparisons across different models. In contrast, traditional methods such as SHAP, PDPs, and counterfactuals are more effective for explaining individual predictions or identifying which features drive decisions. H-XAI combines both, so stakeholders can move between overall model behavior and instance-level explanations, depending on the questions they care about. To ground our proposal, we present two case studies spanning six scenarios: credit risk classification using the German Credit dataset and financial time-series forecasting with stock prices. In both cases, we show how different types of questions posed by applicants, regulators, or data scientists can be addressed using appropriate combinations of methods.
Our contributions are as follows: 1. We introduce Holistic-XAI (H-XAI), a unified formulation that combines black-box ratings with post-hoc XAI techniques to support both individual- and system-level explanation needs (Figure 1(c)); 2. We propose a rating-driven explanation (RDE) workflow that allows hypothesis testing using causal metrics, and supports comparative model evaluation using automatically generated random and biased baselines (Figure 2); 3. We demonstrate how H-XAI supports a range of stakeholder needs through two representative case studies: binary classification and time-series forecasting, and summarize method selection strategies across questions and contexts (Section 4 and Appendix D).
Figure 1: (a) Generalized Causal Diagram; (b) Stakeholder types in the context of Blackbox AI; (c) H-XAI Workflow: End-user receives output from a blackbox AI. Stakeholders may seek related explanations. “Method” shows type (.) and instances from case studies.
2 Background and Related Work
In this section, we describe how we categorized stakeholders by drawing on prior work, and how we selected a subset of these stakeholder types to represent in our case studies in Section 4. We also explain how the questions we chose reflect the explanation needs identified in the literature, and conclude by situating our work within relevant prior work. An extended discussion of related work, including a dedicated section on explainability for time-series models, is available in the appendix (Section A).
2.1 Background: Stakeholders and Their Explanation Needs
We follow the stakeholder taxonomy from (Deshpande and Sharp 2022), which identifies three main groups (Figure 1(b)):
1. Individual stakeholders: End users directly impacted by model predictions. Their questions often focus on personal outcomes and decision justification.
2. Regulatory stakeholders: Policy-makers and auditors concerned with fairness, bias, and legal compliance.
3. Organizational stakeholders: Includes operational roles (e.g., business leads, data scientists) and technical roles (e.g., developers) concerned with performance and reliability.
To structure their explanation needs, we adapt the XAI question taxonomy from (Liao et al. 2021), which organizes nine categories of explanatory questions from interviews with industry practitioners. Though originally domain-agnostic, many of their core questions, such as “Why was this prediction made?”, “What would change the outcome?”, or “Which features matter most?”, align closely with our stakeholder scenarios in credit scoring and forecasting. We pair each question type with a fitting method: SHAP for feature attribution, counterfactuals for what-if queries, and RDE for behavior under perturbations or across subgroups. A full mapping is provided in the appendix (Table D.2).
2.2 Explainability in AI
Explainable AI (XAI) methods aim to clarify models behavior by describing predictions, decision patterns, and potential risks (Arrieta et al. 2020). Toolkits like AIX360 (Arya et al. 2019) offer diverse techniques to support this goal. However, most explanations remain static, built to justify outcomes rather than support interaction or hypothesis testing. Prior work argues that explanation is not a fixed output but a process of exploration (Mueller et al. 2019; Hoffman et al. 2023a). Users often need to ask follow-up questions or test alternate scenarios to build meaningful understanding (Abdul et al. 2018; Chromik and Butz 2021). Existing methods rarely support this kind of iterative reasoning or sensitivity analysis. In H-XAI, we use post-hoc methods like SHAP and PDPs to answer questions about feature influence on predictions, both locally and globally. They, however, are limited when users want to test hypotheses, such as whether a model’s decision would change under a counterfactual scenario, or whether outcomes differ systematically across groups. For such cases, we use RDE that estimate perturbation effects, account for confounding, and provide model comparisons.
2.3 Rating of AI Models
Prior work has introduced rating methods to assess AI models from a third-party perspective, often without access to training data. (Srivastava and Rossi 2020) proposed a framework to quantify bias, particularly gender bias, in machine translation systems (Srivastava and Rossi 2018), using visualizations to communicate results (Bernagozzi et al. 2021b, a). However, these approaches lacked a causal foundation. Recent work introduced causal rating methods to isolate the effects of protected attributes on model outcomes. This was first applied to sentiment analysis systems (Lakkaraju, Srivastava, and Valtorta 2024) and then extended to composite NLP tasks (Lakkaraju et al. 2023) and time-series forecasting models (Lakkaraju et al. 2024, 2025). These methods go beyond correlation by estimating treatment effects and adjusting for confounding. In this paper, we build on that foundation but shift the role of ratings: rather than treating them solely as evaluation metrics, we use them to explain model behavior. By integrating ratings with traditional XAI methods, we allow stakeholders to compare models, test hypotheses, and interpret robustness and bias in a more structured way.
2.4 Understanding Stakeholder Needs in XAI
Understanding stakeholder needs is critical for building explanations that are useful, not just technically accurate (Bhatt et al. 2020). For example, (Bundas et al. 2023) highlights NIST’s emphasis on hierarchical vocabularies: high-level concepts should be broadly accessible, while lower levels serve expert users. Studies show stakeholders prefer interactive explanations. Interviews in (Hoffman et al. 2023b) show a strong preference for tools that allow input manipulation and scenario exploration. RDE supports this by allowing users to perturb inputs and observe causal effects on model behavior (Figure D.3). While current XAI methods focus on local explanations, users often want a broader understanding of model behavior (Mueller et al. 2019). (Suresh et al. 2021) proposes a three-level framework for interpretability needs, including the need to understand model limitations. RDE addresses this by showing how behavior changes under different conditions, supporting users with a structured robustness evaluation. In this paper, we show how H-XAI brings these ideas together, supporting interactive, stakeholder-aligned explanations through a combination of rating and post-hoc XAI methods.
3 Approach
Figure 2: RDE Workflow.
To explain model behavior in ways that support robustness, fairness, and practical decision-making, we propose a hybrid framework called Holistic-XAI (H-XAI). It integrates traditional post-hoc explanation methods with rating-driven explanation (RDE) approach. Unlike approaches that rely solely on post-hoc explanations, H-XAI combines both: traditional XAI methods are well-suited for explaining individual predictions and feature attributions, while RDE supports broader questions about robustness, bias, and perturbation effects. This combination lets stakeholders examine both how individual predictions are made and how the model behaves under different input conditions.
3.1 Rating-Driven Explanation (RDE)
RDE combines three components: (a) a causal model that specifies the structural assumptions between key attributes; (b) a set of metrics to quantify causal and statistical bias; and (c) a workflow that connects user queries to ratings and model comparisons.
Causal Model
We begin by specifying a causal model for the task, which includes assumptions about how attributes influence one another. Figure 1(a) shows the generalized causal diagram used across our case studies. While the causal model captures the underlying assumptions, the diagram is its visual representation, making explicit the directional relationships we test. In this setup, the outcome $O$ (e.g., prediction or residual error) is influenced by a treatment variable $T$ (e.g., credit amount, input perturbation), which may itself be influenced by a protected attribute $Z$ (e.g., gender and age). If $Z$ is a common cause of both $T$ and $O$, then confounding exists and must be accounted for to estimate a valid treatment effect.
This diagram allows us to define target causal estimands such as $E[O\mid do(T)]$ (i.e., the deconfounded distribution after adjustment), which represents the expected model outcome under an intervention on $T$. In practice, causal models are often specified by experts, derived from controlled experiments, or inferred using causal discovery algorithms (Gong et al. 2024; Varando et al. 2024), which themselves rely on statistical associations in the observed data. We begin with a hypothesized causal model that assumes protected attributes act as confounders and explore different scenarios within this structure. Even when based on assumptions, such models provide a structured framework for reasoning under interventions, something that correlational methods often fail to do, especially in the presence of confounding (as illustrated by Simpson’s paradox (Pearl 2022)). For instance, in Section 4.1, we consider the possibility that a loan applicant’s status influences the credit amount requested, an entirely plausible scenario in practice. The outcome $O$ may be either the model prediction $\hat{Y}$ or a derived quantity g(Y^)g($\hat${Y}). For example, in time-series settings, we define $O=|\hat{Y}-Y|$ as the residual error. This lets us ask: “Does a certain input perturbation systematically worsen performance for a particular group?”.
Rating Metrics
While RDE builds on prior rating methods, it repositions their function. Earlier works primarily used ratings to compare model performance across input conditions. RDE instead uses them as part of a structured explanation workflow: formulating a question (e.g., “Does treatment T affect outcome O?”), selecting a baseline, applying a metric, and interpreting the result as a potential explanation. This gives ratings an explanatory role, especially in helping stakeholders reason about model bias, instability, or robustness under different conditions. We use three metrics, each designed to answer a specific kind of question; their complete definitions are provided in Appendix B.
1. Weighted Rejection Score (WRS): WRS quantifies statistical differences in outcomes across protected groups using Student’s $t$-test at multiple confidence levels (e.g., 95%, 75%, 60%). A weighted count of rejections across these thresholds yields the final score. WRS is especially useful in settings where the treatment is uniform (e.g., no variation in intervention across groups), such as in our forecasting use case (Section 4.2).
2. Average Treatment Effect (ATE): ATE estimates the causal effect of a treatment on model outcome, averaged over the population. We use propensity score matching (PSM) for deconfounding to improve estimation where needed (Rosenbaum and Rubin 1983).
3. Deconfounded Impact Estimation (DIE %): DIE % quantifies the degree of confounding by measuring the percentage difference between the unadjusted ATE and the ATE estimated after adjusting for the protected attribute (deconfounding).
These metrics serve different explanatory roles: WRS answers “Are outcomes distributed differently across protected attribute groups?”; ATE: answers “Does a treatment cause an outcome change?”; DIE %: answers “Is the observed effect real or driven by confounding?”
RDE Workflow
Figure 2 shows the end-to-end RDE process. Given a user query (e.g., “Is model performance consistent across demographic groups?”), we: 1. Define a treatment ($T$), outcome ($O$), and protected attribute ($Z$), 2. Select the appropriate metric (e.g., ATE or DIE %), 3. Compute scores for the target / test model, 4. Compare it to two automatically generated reference models: - A random model, which outputs predictions at random, - A biased model, which gives different, often extreme predictions based solely on the value of the protected attribute $Z$. For example, in the context of credit risk assessment, always classifying one group as good risk and another group as bad risk.
This comparison grounds the explanation. A model that behaves closer to the biased baseline may raise fairness concerns; one that matches the random model may be unreliable. These comparisons are not just for evaluation: they help users understand why the model’s behavior is problematic and where. RDE outputs can be numeric scores or visual plots, depending on the user query. For example, Figure D.1 shows a case where RDE highlights age-based confounding in a credit classifier, while Figure D.3 shows how error rates change under missing input perturbations.
3.2 Traditional XAI Methods Used
While RDE provides explanations concerning the robustness of the models at a global level, it is not suitable for local feature attribution or communicating the robustness of the model at a global level. To address these needs, we integrate standard post-hoc explanation methods, which stakeholders may already be familiar with. The complete definitions of each can be found in Appendix C.
PDPs: Show the marginal effect of a feature on predictions by averaging over the dataset. Useful for understanding monotonicity or directionality, but unreliable when features are correlated. In Figure D.1, we use PDPs to examine the effect of credit amount across applicant groups.
SHAP: Estimates the marginal contribution of each feature to a prediction, using additive Shapley values. For time-series models, we use TreeSHAP on surrogate models (Raykar et al. 2023) to reduce computational cost. This makes the explanations more scalable, even for long input sequences. In our case study, we evaluated the surrogate on a small subset (10% of data, 60 samples) and found that its MASE and SMAPE differed from MOMENT’s by only 0.66 and 0.18%, respectively (see Table D.1 in the appendix).
Counterfactuals: Identify minimal input changes needed to flip a prediction (Wachter, Mittelstadt, and Russell 2017). This is useful for questions like, “What should this applicant change to get approved?”
3.3 Holistic-eXplainable AI (H-XAI)
H-XAI brings together traditional post-hoc methods, some of which explain individual predictions (like SHAP and counterfactuals) and others that show general patterns (like PDPs or global SHAP), with RDE, which focuses on robustness by testing how model behavior changes under interventions and comparing it to biased and random baselines. This lets us analyze not just how the model works, but how stable and fair it is under different conditions. This process is often driven by stakeholders themselves, who have hypotheses involving certain variables they want to test. Stakeholders may begin with a general question (“Is the model fair across groups?”), use RDE to detect problems, then drill down into specific examples using SHAP or counterfactuals to understand the source of those issues. As shown in Figure 1(c), this setup allows users to move between different types of explanations depending on the question, without being locked into a single method or interpretability style. We demonstrate this combined framework through two use cases in Section 4, mapping specific stakeholder questions to the methods used to answer them.
Figure 3: (a) Causal diagram for binary classification on the German Credit dataset. Protected attributes include Age, Personal Status, Gender, or their combinations, depending on query. (b) Causal diagram for time-series forecasting (stock prices), adapted from (Lakkaraju et al. 2024).
4 Case Studies
Figure 4: Hypothesis: Does decreasing the credit amount by 50% improve the chances of loan approval? Conclusion: On average, decreasing the requested credit amount by 50% increases the risk by 0.06, meaning applicants will be more likely to be classified as good risk (denoted by class 1). This suggests that asking for a smaller loan may improve the likelihood of loan approval only marginally.
Question Method Used Limitation When to Combine
Do outcomes differ across groups (e.g., gender, age)? WRS Cannot isolate cause of disparity (only statistical difference). Combine with DIE % to check if protected attribute has a confounding or direct effect.
Does a certain input (e.g., credit amount) causally influence the outcome? ATE Does not explain at the individual level; ignores confounding if unadjusted Use DIE% to validate causal effect is not confounded
Is the effect of a feature (e.g., credit amount) on the outcome distorted because it is influenced by a protected attribute (or confounding) (e.g., age)? DIE% Cannot be used when treatment is uniform; needs good causal estimation models; not instance-level Use SHAP to identify where and how confounding influences predictions at the instance level
How does a feature influence predictions overall? PDP Fails when features are correlated; only shows average trends Use ATE for more reliable estimation if confounding is present.
Which features contribute most to a single prediction? SHAP (local) Can be unstable with feature interactions; does not show global behavior Use global SHAP or ATE to see if local pattern is consistent across dataset
Which features matter globally? SHAP (global) Attribution unreliable under correlated inputs; may mis-calculate importance Use RDE to test if feature importance remains stable across subgroups or under input perturbations.
What minimal change would flip a decision? Counterfactual Explanation May suggest unrealistic or invalid inputs; ignores global trends Validate counterfactual’s plausibility using SHAP or PDP (helps identify features that matter to the model).
Does the model behave differently for one group under noisy/missing input? DIE% Needs a causal model. Use SHAP to identify which feature contributions shift across groups post-perturbation.
Is the model sensitive to a certain input change? ATE (perturbation as treatment) Only provides average effect; hides heterogeneity. Use SHAP to see variability of effect across instances.
Does this error pattern resemble biased or random behavior? RDE with biased/random baselines Baselines are synthetic; may not match real misuse cases. Combine with counterfactuals or SHAP to examine edge cases or deviations.
Table 1: Mapping explanation methods to stakeholder questions, limitations, and when to combine them. Combinations show scenarios from our case studies (Section 4). Others are possible depending on task and data context.
In this section, we present two case studies to demonstrate how H-XAI can address the explanation needs of different stakeholders (introduced in Section 2.1). These scenarios are structured around a taxonomy of questions that align with the XAI Question Bank (Liao et al. 2021). The mapping of our scenario questions to the categories of the XAI Question Bank is provided in the appendix (Table D.2). We show six scenarios: three for each stakeholder type in a given task. Three are included here, and the rest are in Appendix D.
4.1 Case Study-1: Binary Classification
In the first case study, we choose binary credit risk classification as it is a well-studied, high-stakes domain where fairness, explanation, and model behavior have real-world consequences (Demajo, Vella, and Dingli 2020; Kłosok, Chlebus et al. 2020; Melsom, Vennerød, and de Lange 2022).
A. Data. German Credit Dataset (Frank 2010) has been widely used in prior work on algorithmic fairness and explainability (Zemel et al. 2013; Martins et al. 2023; Verma and Rubin 2018), making it suitable for examining how different stakeholder questions surface in a binary classification task. It contains 1,000 rows with 20 applicant features, including demographic information (e.g., age, personal status and sex), financial details (e.g., credit amount, duration, checking account status), and other applicant-related attributes which are used to assess loan risk. We selectively analyze subsets of these features depending on the specific explanation or causal question being addressed. The primary outcome variable is binary: $Risk\in\{0,1\}$, where $Risk=1$ indicates a good credit risk (leading to loan approval), and $Risk=0$ indicates a bad credit risk (loan rejection).
B. Causal Setup. For generating RDE, we use the causal diagram shown in Figure 3(a). Specifically, we focus on a smaller set of variables: we treat Credit Amount as the treatment, Age and Personal Status and Sex as protected attributes following (Zemel et al. 2013), and Risk as the outcome. In contrast, when applying XAI methods like SHAP, we do use the full feature set, since those methods are not bound by the same identifiability constraints and are designed to give feature attribution-based explanation rather than estimate causal effects.
C. AI Models Considered. We use two models: a linear, interpretable model (Logistic Regression), and a non-linear ensemble model (Random Forest). They are two of the most popular methods used for credit scoring (Addy et al. 2024). Figure D.2 in the appendix highlights an accuracy versus confounding bias trade-off in Scenario-3. While Random Forest outperforms Logistic Regression in accuracy, it also exhibits higher confounding bias. This reflects the common, but not universal, trade-off between accuracy and interpretability (Atrey et al. 2025).
We present three scenarios in which H-XAI (described in Section 3) is applied to address questions posed by different types of stakeholders: an individual applicant (Figure 5), a regulatory official (Figure D.1 in the appendix), and a data scientist within an operational organization (Figure D.2 in the appendix).
Figure 5: Stakeholder: Jack (applicant; individual); Scenario-1: Bank XYZ uses a Random Forest model to classify applicants as good or bad risk based on features like credit amount, age, personal status, gender, and more. Jack was classified as a bad risk and wants to understand why, and what changes he could make to be considered a good risk and get his loan approved.
We also provide an example of a hypothesis-driven explanation for credit risk classification in Figure 4. The causal setup described for RDE in Section 3.1 also supports hypothesis-driven explanations, allowing users to test claims about model behavior rather than receive fixed outputs, as illustrated in Figure 4. In the appendix, we show how doubling age (Figure E.4) or halving loan duration (Figure E.5) affects predicted credit risk, demonstrating how hypothesis-driven explanations can serve as a standalone method for validating assumptions about model behavior.
Figure 6: Stakeholder: Suresh (Regulatory official; Regulatory body). Scenario-4: Suresh uses a stock forecasting tool developed by a third-party company to guide her investment decisions. He notices that the tool performs poorly on certain days and suspects it may be less reliable for some companies. To decide whether to continue using it, he wants to understand how robust the tool is.
Figure 7: Stakeholder: Jane (Retail Investor; Individual). Scenario-5: Jane uses a stock forecasting tool developed by a third-party company to guide her investment decisions. She notices that the tool performs poorly on certain days and suspects it may be less reliable for some companies. To decide whether to continue using it, she wants to understand how robust the tool is.
4.2 Case Study-2: Time-Series Forecasting
In the second case study, we focus on stock price forecasting, a task that is high-impact in financial decision-making and widely used in algorithmic trading, portfolio management, and investment analytics.
A. Data. We collected daily stock prices from Yahoo! Finance for six companies across different industries: Meta (META) and Google (GOO) in social technology, Pfizer (PFE) and Merck (MRK) in pharmaceuticals, and Wells Fargo (WFC) and Citigroup (C) in financial services, as closing stock prices is the most commonly used data source for time-series forecasting (Lin and Marques 2024; Lakkaraju et al. 2024). The dataset covers from March 28, 2023, to April 22, 2024; data up to March 22, 2024, was used to forecast the following month.
B. Causal Setup. To generate RDE for time-series forecasting, we adopt the preprocessing approach from (Lakkaraju et al. 2024), where the difference between the 20 predicted stock prices (based on the previous 80 time steps) and their corresponding ground truth values is used to compute a sequence of residuals. While other metrics like SMAPE or MASE are available, we use residuals because they give the raw prediction error at each point, rather than at an aggregate level. To capture worst-case performance, we take the maximum residual within each prediction window as the outcome of interest. For simplicity, we refer to this as the residual throughout the paper. While predicted stock prices may vary across companies, our focus is on whether residuals, i.e., model errors, also vary systematically, which may indicate unreliability. This residual serves as the basis for analyzing how protected attributes or perturbations affect model behavior. Figure 3(b) shows the causal diagram used for this analysis. We treat Perturbation (e.g., introducing missing values or zero-imputation, as in Figure D.3) as the treatment, Company as the sensitive attribute, and Residual as the outcome. This causal setup is used exclusively for explanation questions that require RDE.
C. AI Models Considered. Our goal is not to benchmark forecasting performance but to analyze and explain model behavior. So rather than selecting the most competitive models, we chose two representative ones from the models used in (Lakkaraju et al. 2024): MOMENT, an open-source, time-series-specific foundation model, and Gemini, a general-purpose foundation model, both of which are easily accessible. We also use ARIMA, a statistical time-series forecasting model, as an additional baseline.
We present three scenarios in H-XAI and apply them to address questions posed by different types of stakeholders: a regulatory official (Figure 6), a retail investor (Figure 7), and a data scientist within an operational organization (Figure D.3 in the appendix).
4.3 When to Use: Mapping Questions to Methods
A key goal of H-XAI is to help users in choosing the right tool for the right question, not only to apply methods indiscriminately. Table 1 maps stakeholder questions to fitting methods. We summarize what each reveals, their limitations, and when they require complementary methods. This is where H-XAI matters: it lets users pose questions, pick methods, identify problems, and cross-validate them using complementary methods. No single method is sufficient; the strength lies in the combination.
5 Discussion and Conclusion
While our case studies focus on credit risk classification and stock price forecasting, H-XAI is task-agnostic and extends beyond these domains. It emphasizes selecting explanation techniques based on stakeholder roles and questions, rather than defaulting to a single method. In classification tasks, RDE may estimate the causal impact of applicant attributes on model decisions, while in forecasting, it can quantify how input perturbations affect performance. Traditional XAI methods complement RDE by providing local, instance-level explanations where causal analysis may not apply.
As shown in Figure 2, RDE is not just a metric computation step, but a structured reasoning process: constructing baselines, defining a causal model, and selecting appropriate causality-based metrics to answer the question. This structure enables H-XAI to adapt to different stakeholder needs, whether the goal is to understand robustness, quantify bias, or interpret individual predictions, making it suitable for critical domains like healthcare and finance. We also recognize that stakeholders may not always know how to define a causal model. To address this, RDE uses a generalized causal model by default, capturing common structural assumptions across tasks. Whether the analysis targets confounding bias (e.g., DIE %) or statistical bias (e.g., WRS) depends on how the question is framed, as demonstrated in Q1 of Figure 5 versus Q1 of Figure D.3. We do not treat interpretability as a fixed property of a method. Each explanation is shaped by the stakeholder’s question and the context.
A key takeaway from our work is that explanation is an iterative process, not a static output. No single method is sufficient on its own; different questions demand different tools. We show that by selecting and combining methods, whether RDE, counterfactuals, or feature attribution, we can shift the role of explanation from justifying predictions to supporting meaningful interaction, allowing stakeholders to test hypotheses and refine their understanding. While a promising step toward a unified approach, this work opens several avenues for future research, including expanded case studies, additional methods, and conducting user studies with diverse stakeholders.
Appendix
Appendix A Related Work
A.1 Explainability in AI
Explainable AI (XAI) is a set of processes and methods that are used to describe AI models, by characterizing their accuracy, transparency, and outcomes (Arrieta et al. 2020). AIX360 (Arya et al. 2019) introduces an open-source toolkit for explainability, featuring a range of SOTA techniques to guide users in selecting appropriate explanation methods based on their needs. While this toolkit is very valuable, the explanations do not allow users to interact with them. (Mueller et al. 2019) consider an effective explanation to be an interaction and not a property of statements. (Abdul et al. 2018) also acknowledges that a single explanation rarely results in a desired level of understanding. Explanation is an exploratory activity and never a ”one-off” (Hoffman et al. 2023a). (Chromik and Butz 2021) categorizes prior XAI work into different types of interaction based on the concepts introduced by (Hornbæk and Oulasvirta 2017). The paper highlights that prior research has been limited in facilitating natural, iterative conversations about AI behavior, an essential aspect for achieving transparency and scrutability in XAI systems. Additionally, such exploratory dialogues allow users to explore how changes in input lead to changes in the AI prediction (Chromik and Butz 2021). Through H-XAI, we address a series of user questions, where each question builds upon the previous one. This iterative process culminates in suggesting potential steps that the user can take to achieve the desired model outcome.
Feature importance methods such as SHapley Additive exPlanations (SHAP) and counterfactual explanations are the most widely-used explanation methods. The former is sensitive to feature correlations and the latter only explains a single outcome instead of the entire model (Alfeo et al. 2023). Relying on People benefit more from having local and global explanations that are integrated, yet the majority of XAI methods have focused on the delivery of local explanations (Hoffman et al. 2023b). We employ correlation-based XAI methods to provide high-level explanations of model decisions, while for more in-depth user questions that require understanding the underlying causal structure, we transition to causality-based explanations to offer deeper insights into the model’s behavior.
A.2 Explainability in Time-Series Forecasting Models (TSFMs)
Traditional XAI methods for TSFMs rely on feature attribution and visualization methods. Model-agnostic methods such as SHAP are widely used (Mokhtari, Higdon, and Başar 2019; Schlegel et al. 2019; Zhang et al. 2022), along with partial dependence plots (Greenwell 2017) to estimate marginal contributions of input variables and forecast outputs. These methods have been adapted to consider time-dependent inputs, but they are often limited in capturing causality or robustness under input perturbation (Lakkaraju et al. 2025). In deep learning-based TSFMs, attention mechanisms have also been used to show influential time steps or variables in multivariate forecasting tasks (Abbasimehr and Paki 2022; Li et al. 2019; Shih, Sun, and Lee 2019). These were used to picture instance-level temporal interpretability. However, such mechanisms are typically embedded within the models and therefore may not be reliable indicators to measure causal influence. Though these methods have been useful for prediction interpretability, they still tend to disproportionately focus on either local or global explanations as opposed to both. SHAP often provides local insights into single predictions. PDPs, for instance, offer aggregated trends across a dataset. In TSFMs, it is often necessary to combine both perspectives, i.e., understanding how individual time points contribute to certain forecasts (local), and how these patterns also generalize across entire data (global). Existing approaches do not always consistently support both forms of explanation in an integrated or interactive manner, especially under stakeholder-specific constraints. In recent years, there has been a growing shift in the literature toward methods that not only explain model predictions but also empower stakeholders at different levels to interrogate model behavior through hypothesis testing and counterfactual reasoning (Çelik, İcan, and Bulut 2023; Tasin et al. 2023). Motivated by this trend, our work supports stakeholder-aligned explanation for TSFMs (Section 4.2) across two key dimensions: accuracy and robustnes, which includes performance under varying input conditions as well as bias.
Appendix B RDE Metrics
Weighted Rejection Score (WRS): WRS quantifies statistical bias across a protected attribute $z_{i}\in Z$ by comparing outcome distributions $(O\mid z_{i})$ across all $\binom{m}{2}$ group pairs using Student’s t-test (Student 1908). For each pair, the t-value $t_{z_{i}}$ is compared to the critical value $t_{\text{crit}}$ for a given confidence interval (CI). A rejection indicates a significant difference in outcomes. WRS is computed as a weighted sum of rejections across CIs, where $x_{i}$ is the number of rejections and $w_{i}$ is the assigned weight (e.g., 1, 0.8, 0.6 for 95%, 75%, and 60% CIs). Formally, $WRS=\sum_{i}w_{i}\cdot x_{i}$.
We compute WRS when the treatment is applied uniformly across all protected attribute groups, making DIE % inappropriate in such cases (as in Section 4.2).
Average Treatment Effect (ATE): Average Treatment Effect (ATE) (Rubin 1974; Splawa-Neyman, Dabrowska, and Speed 1990) measures the average difference in outcomes between the treated ($T=i$) and untreated ($T=0$) units, quantifying the impact of the treatment on the AI model’s outcome. Unlike local counterfactuals, ATE quantifies the average effect of a treatment on model predictions across the dataset, making it well-suited for answering stakeholder questions about overall effect of an attribute on the outcome. Formally, ATE=[|E[O|do(T=i)]E[O|do(T=0)]|]ATE=[|E[O|do(T=i)]-E[O|do(T=0)]|]
Deconfounded Impact Estimation (DIE %): DIE % measures how much confounders ($Z$) distorts the relationship between a treatment ($T$) and an outcome ($O$), leading to confounding bias. We use Propensity Score Matching (PSM) to calculate DIE % when dealing with categorical treatments, but extend it to continuous treatments using G-computation (Robins 1986). Instead of linear models, we use a SuperLearner (Van der Laan, Polley, and Hubbard 2007), an ensemble of non-linear models to estimate counterfactual outcomes. DIE % evaluates the difference between the ATE computed without adjustment ($ATE_{0}$) and the ATE after applying G-computation ($ATE_{g}$). Formally, $DIE\%=\left[||ATE_{o}|-|ATE_{g}||\right]\times 100$
Here is how the rating metrics are calculated.
  • Algorithm 1 computes the Weighted Rejection Score (WRS) to measure statistical bias by analyzing how sensitive the model’s outcomes are to variations in sensitive attributes.
  • Algorithm 2 calculates Deconfounding Impact Estimation (DIE %) using PSM to assess confounding bias by measuring the effect of the confounder on model outcomes before and after deconfounding.
  • Algorithm 3 calculates ATE by evaluating the difference in the model’s outcomes for treated data (perturbation applied) and control data (no perturbation) to evaluate the impact of the treatment on the outcome. The ATE is computed after applying PSM or G-computation to mitigate the confounding effects, hence the use of the do(.) operator in the ATE formulation.
Algorithm 1 WeightedRejectionScore
Purpose: Calculate the weighted sum of null-hypothesis rejections for dataset $d_{j}$ and model $s$ using confidence intervals $CI_{k}$ and weights $w_{k}$.
Input:
$d$: Dataset for a specific perturbation;
CICI: Confidence intervals (e.g., 95%, 70%, 60%);
$s$: AI model from the set of test models $S$;
$W$: Weights for different CIs (e.g., 1, 0.8, 0.6)
Output: $\psi$: Weighted rejection score
[…]
2:for each $(ci_{i},w_{i})$ in $(CI,W)$ do
3:  for each pair $(z_{a},z_{b})$ in classes of $Z$ do
4:   $(t,pval,dof)\leftarrow\text{T-Test}(z_{a},z_{b})$
5:   $t_{crit}\leftarrow\text{LookUp}(ci_{i},dof)$
6:   if $t>t_{crit}$ then
7:     $\psi\leftarrow\psi+w_{i}$
8:   end if
9:  end for
10:end for
11:return $\psi$
Algorithm 2 ComputeDIEScore
Purpose: Calculate the Deconfounding Impact Estimation (DIE) using Propensity Score Matching (PSM).
Input:
$s$: AI model,
$d$: Dataset corresponding to a perturbation,
$p$: A perturbation other than $p_{0}$ (control or no perturbation)
Output: $\psi$: DIE score
[…]
4:return $\psi$
Algorithm 3 ComputeATEScore
Purpose: Calculate the Average Treatment Effect (ATE).
Input:
$s$: AI model,
$d$: Dataset,
$p$: Treatment perturbation,
$p_{0}$: Control (unperturbed baseline)
Output: $\psi$: ATE score
[…]
2:return $\psi$
Appendix C Traditional XAI Methods Used
In this paper, we use widely recognized and established post-hoc XAI approaches in Finance domain (Martins et al. 2023) (as both the case studies illustrated in Section 4 are related to finance) such as SHAP, PDPs, and counterfactual explanations in this paper. While more advanced methods have emerged, in our paper, our focus remained on how these well-understood techniques can effectively answer specific questions. Each can be easily substituted with more sophisticated alternatives, such as DeepLIFT (Li et al. 2021) or DeepSHAP (Chen, Lundberg, and Lee 2022), depending on the complexity and needs of the application. We only focus on post-hoc explanation methods as they can be easily applied to a trained model by anyone.
Partial Dependence Plots (PDPs).
We use PDPs to visualize how model predictions vary with a specific feature $x_{j}$, averaged across the dataset (Friedman 2001). PDPs help interpret the model’s behavior by isolating the marginal effect of a single feature while averaging out the influence of all other features. This makes them particularly useful for understanding global model behavior in response to changes in individual variables. The partial dependence function for feature $x_{j}$ is formally defined as: $\text{PDP}(x_{j})=\frac{1}{n}\sum_{i=1}^{n}\hat{f}(x_{j},x_{k}^{(i)})$.
Here, $\hat{f}$ denotes the model’s prediction function, $x_{k}^{(i)}$ represents all features except $x_{j}$ for the $i$-th instance, and the result is the average predicted outcome when $x_{j}$ is fixed and all other features are held at their observed values. We use PDPs when the goal is to understand overall trends in how a feature influences predictions. However, ATE provides a more precise and nuanced estimate, especially in the presence of confounders, and can sometimes replace PDP when the aim is to quantify group-level differences more reliably (for e.g., Q2 in Figure D.1).
SHapley Additive exPlanations (SHAP).
We use SHAP (Lundberg and Lee 2017) to interpret how individual features contribute to a model’s prediction. SHAP assigns each feature a value $\phi_{j}$ that reflects its marginal contribution to the prediction, averaged over all possible subsets of features. The explanation is given by an additive model: $\hat{f}(x)=\phi_{0}+\sum_{j=1}^{k}\phi_{j}z^{\prime}_{j}$.
where $\phi_{0}$ is the expected prediction, and $z^{\prime}_{j}\in\{0,1\}$ indicates whether feature $j$ is included in a given coalition. SHAP can be applied locally to explain single predictions, or globally by aggregating values across the dataset.
In Section 4.2, computing SHAP values directly on large models (e.g., foundation models) requires $\mathcal{O}(2^{M})$ model evaluations per instance, where $M$ is the number of input features (e.g., time steps), making it impractical for high-dimensional time-series data. Following TsSHAP (Raykar et al. 2023), we train a tree-based surrogate model and apply TreeSHAP (Lundberg, Erion, and Lee 2018), reducing the explanation cost to $\mathcal{O}(TLD^{2})$, where $T$, $L$, and $D$ are the number of trees, maximum leaves, and depth of the trees, respectively. While (Raykar et al. 2023) uses the surrogate to generate one-step-ahead forecasts recursively to cover the full horizon, we focus only on explaining the first step ahead in our analysis.
Counterfactual Explanation.
Counterfactual explanation (Wachter, Mittelstadt, and Russell 2017) is a local XAI method that describes the smallest change made to one or more input features ($x^{\text{cf}}$) to change the prediction to a desired output ($y^{\text{target}}$).
$f(x^{\text{cf}})=y^{\text{target}},\quad\text{subject to}\quad\min\|x^{\text{cf}}-x\|$.
Appendix D Additional Scenarios from Case Studies
Model SMAPE (%) MASE
Surrogate 7.88 4.77
MOMENT 8.06 5.45
Table D.1: MASE and SMAPE scores on a 10% evaluation subset (60 samples) for the MOMENT model and its surrogate. The small differences indicate that the surrogate closely approximates MOMENT’s predictions for explanation purposes.
Scenario Question in Paper Mapped Question from XAI Question Bank
Case Study 1: Binary Classification (Credit Risk)
Individual (Fig. 5) Q1: On my data instance, I have observed the AI model used by the bank to be biased… How can I rate this model for expected behavior? Performance: How reliable (fairness) are the predictions?
Individual (Fig. 5) Q2: How did each factor… contribute to my loan rejection? Why: What feature(s) of this instance determine the system’s prediction of it
Individual (Fig. 5) Q3: What is the smallest change I could make to my loan duration… to get my loan approved? How to be that: What is the minimum change required for this instance to get a different prediction?
Individual (Fig. 5) Q4: What is the smallest change I could make to my checking account balance to get my loan approved? How to be that: What is the minimum change required for this instance to get a different prediction?
Regulatory (Fig. D.1) Q1: An applicant reported that the model… is biased w.r.t personal status and gender. How can I rate this model for expected behavior? Performance: How reliable (fairness) are the predictions?
Regulatory (Fig. D.1) Q2: Overall, how does the predicted risk change with the applicant’s personal status and gender? How: How does the system make predictions?
Regulatory (Fig. D.1) Q3: How does the model treat female applicants compared to male applicants with varying personal/marital status differently? How: How does the system make predictions? (Specifically, its behavior across subgroups)
Data Scientist (Fig. D.2) Q1: How can I rate the models that are available in my company for accuracy? Performance: How accurate are the predictions?
Data Scientist (Fig. D.2) Q2: How can I rate the models for their bias w.r.t personal status, gender, and age? Performance: How reliable (fairness) are the predictions?
Data Scientist (Fig. D.2) Q3: On average, how does each feature influence the predicted credit risk in the logistic regression model? How: How does the system make predictions?
Case Study 2: Time-Series Forecasting
Retail Investor (Fig. 7) Q1: What in the recent price trend suggests that Google’s stock will drop tomorrow? Why: What features of this instance determine the system’s prediction of it?
Retail Investor (Fig. 7) Q2: What would the forecast be if the stock price at the most influential past time step were \$100 lower? What If: What would the system predict if this instance changes to…?
Retail Investor (Fig. 7) Q3: Does the forecasting model exhibit similar prediction error for another company with similar trends? Why: Why are [instance A and B] given the same prediction?
Retail Investor (Fig. 7) Q4: I observed a local disparity in prediction error between two companies… How can I rate the model’s overall performance by assessing whether such disparities occur more broadly? Performance: How reliable (fairness) are the predictions?
Regulatory (Fig. 6) Q1: A user reported that the forecasting tool… relies heavily on older stock prices. Does the model overall, depend more on long-term patterns than recent ones? How: How does it weigh different features?
Regulatory (Fig. 6) Q2: Given this reliance on long-term patterns, can the model still be trusted to consistently predict the correct direction of stock price movement across companies? Performance: What are the limitations of the system? (specifically, its reliability)
Data Scientist (Fig. D.3) Q1: How can I rate the models that are available… for their accuracy and bias w.r.t Company? Performance: How accurate / reliable (fairness) are the predictions?
Data Scientist (Fig. D.3) Q2: How can I rate these models for their accuracy and bias w.r.t Company when there are missing values in the data? Performance: How reliable (fairness and sensitivity to input perturbations (missing values)) are the predictions?
Data Scientist (Fig. D.3) Q3: After replacing the missing values with zero, how can I rate the models for their accuracy and bias w.r.t Company? Performance: How reliable (fairness and sensitivity to input perturbations (zeros)) are the predictions?
Table D.2: Mapping of all scenario questions to the XAI Question Bank from (Liao et al. 2021). The highlighted questions in the table are specifically focused on model fairness and sensitivity to perturbations. While the question bank lacks explicit questions for these topics, we have mapped our questions to the closest available ones, highlighting a need for the original taxonomy to be extended to include these topics.
Figure D.1: Stakeholder: Maya (Regulatory official; Regulatory body). Scenario-2: An applicant reports that the AI model used by Bank XYZ is biased with respect to personal status and gender, and suspects their loan was rejected for this reason. Maya, a regulator from the ABC Bureau, begins an investigation to determine whether this bias is present.
Figure D.2: Stakeholder: James (Data Scientist; Operational Organization). Scenario-3: Given a set of candidate models, James sets out to evaluate which one performs best on the client’s credit risk data. He plans to assess both accuracy and fairness before selecting a model for deployment; Interestingly, the global SHAP summary (Q2) is closely aligned with the Jack’s SHAP values shown in Figure 5!
Figure D.3: Stakeholder: Amanda (Data scientist; Operational organization). Scenario-6: Amanda is evaluating several time-series forecasting models for deployment. She wants to compare their accuracy and bias to choose one that best fits her team’s priorities.
Figures D.1,D.2,D.3 show additional scenarios from case studies illustrated in Section 4.
Table D.2 shows how the questions from the paper’s case studies align with an existing XAI taxonomy (Liao et al. 2021). The table highlights specific questions on model fairness and sensitivity to perturbations, which lack explicit counterparts in the question bank.
Appendix E Hypothesis-driven Explanation
We define hypothesis-driven explanations as a structured process in which users test specific assumptions or causal questions about model behavior, rather than passively receiving fixed outputs. While this is not exactly what Miller envisioned, it aligns with the direction of his Evaluative AI framework (Miller 2023). In our formulation, such hypotheses are operationalized through causal models, allowing users to quantify effects and systematically examine how the model behaves. Figure 4, Figure E.4, and Figure E.5 show different examples in the context of credit risk assessment using German credit dataset which was previously used to demonstrate H-XAI in Section 4.1. Age, Personal status, and Gender are assumed to act as confounders for both credit amount and loan duration in months.
Figure E.4: Hypothesis: Does doubling applicants’ age, reduce their chances of being approved for a loan? Conclusion: On average, doubling a person’s age led to a noticeable drop in the predicted creditworthiness (risk score decreased by 0.35). This suggests that the model tends to classify older applicants as higher risk, which could be a sign of age-related bias.
Figure E.5: Hypothesis: Does decreasing the loan duration by half improve the chances of loan approval? Conclusion: Reducing the loan duration by half led to a slight increase in predicted creditworthiness (risk score increased by 0.097). This indicates that the model tends to view shorter loan durations more favorably, which may slightly help an applicant get approved.
Appendix F Source Code for Data Processing
1# Convert data from Yahoo! Finance to sliding window format.
2def sliding_window(data, window_size, company):
3 sequences = []
4 for i in range(len(data) - window_size):
5 seq = data[i:(i + window_size + 1)].tolist()
6 sequences.append([company] + seq)
7 return pd.DataFrame(sequences)
8
9# Perturbations:
10# Drop-to-zero: Every 80th stock price in the numerical data will be turned into zero.
11def drop_to_zero(df, col):
12
13 new_df = df.copy()
14 new_df.loc[new_df.index % 80 == 0, col] = 0
15
16 return new_df
17
18# Missing values: Every 80th stock price in the numerical data will be NaN’.
19def missing_values(df, col):
20
21 new_df = df.copy()
22 new_df.loc[new_df.index % 80 == 0, col] = float(’nan’)
23
24 return new_df
Listing 1: Preprocessing script used to prepare the German Credit dataset.
1# Reference: https://www.kaggle.com/code/pahulpreet/german-credit-risk-analysis-beginner-s-guide/notebook
2
3df = pd.read_csv("http://archive.ics.uci.edu/ml/machine-learning-databases/statlog/german/german.data", sep=" ", header=None)
4headers = ["Status of existing checking account", "Duration in month", "Credit history",
5 "Purpose", "Credit amount", "Savings account/bonds", "Present employment since",
6 "Installment rate in percentage of disposable income", "Personal status and sex",
7 "Other debtors / guarantors", "Present residence since", "Property", "Age in years",
8 "Other installment plans", "Housing", "Number of existing credits at this bank",
9 "Job", "Number of people being liable to provide maintenance for", "Telephone",
10 "foreign worker", "Cost Matrix(Risk)"]
11df.columns = headers
12
13df["Status of existing checking account"] = df["Status of existing checking account"].map({
14 ’A14’: "no checking account", ’A11’: "<0 DM", ’A12’: "0 <= <200 DM", ’A13’: ">= 200 DM"
15})
16
17df["Credit history"] = df["Credit history"].map({
18 "A34": "critical account", "A33": "delay in paying off", "A32": "existing credits paid back duly till now",
19 "A31": "all credits at this bank paid back duly", "A30": "no credits taken"
20})
21
22df["Purpose"] = df["Purpose"].map({
23 "A40": "car (new)", "A41": "car (used)", "A42": "furniture/equipment", "A43": "radio/television",
24 "A44": "domestic appliances", "A45": "repairs", "A46": "education", "A47": "vacation",
25 "A48": "retraining", "A49": "business", "A410": "others"
26})
27
28df["Savings account/bonds"] = df["Savings account/bonds"].map({
29 "A65": "no savings account", "A61": "<100 DM", "A62": "100 <= <500 DM",
30 "A63": "500 <= <1000 DM", "A64": ">= 1000 DM"
31})
32
33df["Present employment since"] = df["Present employment since"].map({
34 ’A75’: ">=7 years", ’A74’: "4<= <7 years", ’A73’: "1<= < 4 years", ’A72’: "<1 years", ’A71’: "unemployed"
35})
36
37df["Personal status and sex"] = df["Personal status and sex"].map({
38 ’A95’: "female:single", ’A94’: "male:married/widowed", ’A93’: "male:single",
39 ’A92’: "female:divorced/separated/married", ’A91’: "male:divorced/separated"
40})
41
42df["Other debtors / guarantors"] = df["Other debtors / guarantors"].map({
43 ’A101’: "none", ’A102’: "co-applicant", ’A103’: "guarantor"
44})
45
46df["Property"] = df["Property"].map({
47 ’A121’: "real estate", ’A122’: "savings agreement/life insurance",
48 ’A123’: "car or other", ’A124’: "unknown / no property"
49})
50
51df["Other installment plans"] = df["Other installment plans"].map({
52 ’A143’: "none", ’A142’: "store", ’A141’: "bank"
53})
54
55df["Housing"] = df["Housing"].map({
56 ’A153’: "for free", ’A152’: "own", ’A151’: "rent"
57})
58
59df["Job"] = df["Job"].map({
60 ’A174’: "management/ highly qualified employee", ’A173’: "skilled employee / official",
61 ’A172’: "unskilled - resident", ’A171’: "unemployed/ unskilled - non-resident"
62})
63
64df["Telephone"] = df["Telephone"].map({
65 ’A192’: "yes", ’A191’: "none"
66})
67
68df["foreign worker"] = df["foreign worker"].map({
69 ’A201’: "yes", ’A202’: "no"
70})
71
72df["Cost Matrix(Risk)"] = df["Cost Matrix(Risk)"].map({
73 1: "Good Risk", 2: "Bad Risk"
74})
75
76# Label encoding categorical variables
77from sklearn.preprocessing import LabelEncoder
78label_encoder = LabelEncoder()
79categorical_columns = df.columns[~df.columns.isin([’Age in years’, ’Duration in month’, ’Credit amount’])]
80for column in categorical_columns:
81 df[column] = label_encoder.fit_transform(df[column])
82 mapping = dict(zip(range(len(label_encoder.classes_)), label_encoder.classes_))
[…]
Appendix G Additional Implementation Details
All Forecasting Models were executed on Colab notebooks utilizing the L4 GPU available through Colab Pro, which offers 22.5 GB of GPU RAM. Additional details regarding the models, such as the inference times and other architectural details, can be found in Table G.3.
Hyperparameters set
  • MOMENT: head_dropout: 0.1, weight_decay: 0, freeze_encoder: True, freeze_embedder: True, freeze_head: False
  • Gemini: Temperature: 0. Rest of the parameters were default.
Model Mode Size Purpose & Architecture Inf. Time (sec/sample)
Gemini 1.5 Flash Multi-modal 32B* General-purpose, Decoder-only 1.6
MOMENT-large Uni-modal 385M Time-series specific, Encoder-only 0.315
Table G.3: Comparison of Gemini and MOMENT architectures.