### v2

# Holistic Explainable AI (H-XAI): Extending Transparency Beyond Developers in AI-Driven Decision Making

Abstract.
As AI systems increasingly mediate decisions in domains such as credit scoring and financial forecasting, their lack of transparency and bias raise critical concerns for fairness and public trust. Existing explainable AI (XAI) approaches largely serve developers, focusing on model justification rather than the needs of affected users or regulators. We introduce Holistic eXplainable AI (H-XAI), a framework that integrates causality-based rating methods with post-hoc explanation techniques to support transparent, stakeholder-aligned evaluation of AI systems deployed in online decision contexts. H-XAI treats explanation as an interactive, hypothesis-driven process, allowing users, auditors, and organizations to ask questions, test hypotheses, and compare model behavior against automatically generated random and biased baselines. By combining global and instance-level explanations, H-XAI helps communicate model bias and instability that shape everyday digital decisions. Through case studies in credit risk assessment and stock price prediction, we show how H-XAI extends explainability beyond developers toward responsible and inclusive AI practices that strengthen accountability in sociotechnical systems.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
1. Introduction
Explainable AI (XAI) has long promised to help people understand how AI systems behave. Yet despite the variety of methods available, most existing tools are still designed for developers or model builders, not for the wider range of stakeholders who interact with or are affected by AI systems (Mueller et al., 2019; Bhatt et al., 2020; Deshpande and Sharp, 2022). Many such systems now appear in settings where decisions are delivered through online interfaces or web-based services, which further broadens who may need to interpret a system’s behavior. Figure 1(b) illustrates three key stakeholder groups: individual users (e.g., applicants using online decision tools), operational organizations (e.g., data science teams), and regulatory bodies.
Consider the task of credit risk assessment. An applicant might ask, “Would my loan have been approved if I had requested a smaller amount?” A data scientist may wonder, “Which applicant features drive approval decisions on average?” A regulator might want to know whether a model systematically favors certain demographic groups. These questions differ in scope and cannot be handled by a single explanation method. Current XAI tools typically focus on justifying a model’s output rather than helping stakeholders examine it from different angles. They rarely support moving between alternative hypotheses (Hoffman et al., 2023b), testing counterfactuals (Miller, 2023), or identifying when a model behaves unreliably (Suresh et al., 2021). They also tend to produce a fixed explanation, even though practical use often requires several rounds of inspection (Hoffman et al., 2023a).
Past work on black-box model rating offers one way to study model behavior without access to training data. (Srivastava and Rossi, 2018) introduced a two-step procedure to quantify bias in AI services, followed by visualization tools for exploring accuracy-fairness trade-offs (Bernagozzi et al., 2021b). These ideas have since been applied to chatbots (Srivastava et al., 2020) and search engines (Tian et al., 2023). While such ratings communicate how a system behaves “in the wild,” they do not answer the questions stakeholders often ask about why a particular decision changed or how sensitive the model is to certain inputs.
Causal fairness methods go a step further by isolating the influence of protected attributes on predictions (Carey and Wu, 2022). This mirrors the kinds of “what if this were different?” questions people tend to ask. Building on this line of work, (Lakkaraju, 2022; Srivastava et al., 2024) proposed a causal rating framework applied across domains ranging from sentiment analysis (Lakkaraju et al., 2024b) to time-series forecasting (Lakkaraju et al., 2024a, 2025). These methods quantify statistical and confounding bias and evaluate model stability under perturbations. However, they do not provide the instance-level explanations often needed in operational contexts.
In this paper, we propose a framework, Holistic XAI (H-XAI), that brings together traditional XAI techniques and rating-driven explanations (RDE). RDE is helpful for questions involving model instability, bias, and comparisons across models; traditional methods such as SHAP, PDPs, and counterfactuals address local explanations and feature-level reasoning. H-XAI lets stakeholders move between these viewpoints depending on the question at hand. To illustrate how this works, we present two case studies covering six concrete scenarios: credit risk classification using the German Credit dataset and financial time-series forecasting with stock prices. Across these scenarios, we show how the combination of methods supports the kinds of questions raised by applicants, regulators, and data scientists.
Our contributions are as follows:
  1. (1)
    We introduce Holistic XAI (H-XAI), a formulation combining black-box ratings with post-hoc XAI techniques for both instance- and system-level analysis (Figure 1(c));
  2. (2)
    We develop a rating-driven explanation (RDE) workflow that supports hypothesis testing with causal metrics and comparative evaluation using automatically generated random and biased baselines (Figure 2);
  3. (3)
    We demonstrate how H-XAI addresses different stakeholder questions across binary classification and time-series forecasting (Section 4), and discuss practical method-selection strategies (Table 1).
(a) Causal Graph
(b) AI Stakeholders
(c) H-XAI Workflow
Figure 1. (a) Generalized Causal Graph; (b) Stakeholder types in the context of Blackbox AI; (c) H-XAI Workflow: End-user receives output from a blackbox AI. Stakeholders may seek related explanations. “Method” shows type (.) and instances from case studies.
2. Background and Related Work
In this section, we describe how we categorized stakeholders by drawing on prior work, and how we selected a subset of these stakeholder types to represent in our case studies in Section 4. We also explain how the questions we chose reflect the explanation needs identified in the literature, and conclude by situating our work within relevant prior work.
2.1. Background: Stakeholders and Their Explanation Needs
We follow the stakeholder taxonomy from (Deshpande and Sharp, 2022), which identifies three main groups (Figure 1(b)):
1. Individual stakeholders: End users directly impacted by model predictions. Their questions often focus on personal outcomes and decision justification.
2. Regulatory stakeholders: Policy-makers and auditors concerned with fairness, bias, and legal compliance.
3. Organizational stakeholders: Includes operational roles (e.g., business leads, data scientists) and technical roles (e.g., developers) concerned with performance and reliability.
To structure their explanation needs, we adapt the XAI question taxonomy from (Liao et al., 2021), which organizes nine categories of explanatory questions from interviews with industry practitioners. Though originally domain-agnostic, many of their core questions, such as “Why was this prediction made?”, “What would change the outcome?”, or “Which features matter most?”, align closely with our stakeholder scenarios in credit scoring and forecasting. We pair each question type with a fitting method: SHAP for feature attribution, counterfactuals for what-if queries, and RDE for behavior under perturbations or across subgroups. A full mapping is provided in the appendix (Table A.2).
2.2. Explainability in AI
Explainable AI (XAI) methods aim to clarify models behavior by describing predictions, decision patterns, and potential risks (Arrieta et al., 2020). Toolkits like AIX360 (Arya et al., 2019) offer diverse techniques to support this goal. However, most explanations remain static, built to justify outcomes rather than support interaction or hypothesis testing. Prior work argues that explanation is not a fixed output but a process of exploration (Mueller et al., 2019; Hoffman et al., 2023a). Users often need to ask follow-up questions or test alternate scenarios to build meaningful understanding (Abdul et al., 2018; Chromik and Butz, 2021). Existing methods rarely support this kind of iterative reasoning or sensitivity analysis. In H-XAI, we use post-hoc methods like SHAP and PDPs to answer questions about feature influence on predictions, both locally and globally. They, however, are limited when users want to test hypotheses, such as whether a model’s decision would change under a counterfactual scenario, or whether outcomes differ systematically across groups. For such cases, we use RDE that estimate perturbation effects, account for confounding, and provide model comparisons.
2.3. Rating of AI Models
Prior work has introduced rating methods to assess AI models from a third-party perspective, often without access to training data. (Srivastava and Rossi, 2020) proposed a framework to quantify bias, particularly gender bias, in machine translation systems (Srivastava and Rossi, 2018), using visualizations to communicate results (Bernagozzi et al., 2021b, a). However, these approaches lacked a causal foundation. Recent work introduced causal rating methods to isolate the effects of protected attributes on model outcomes. This was first applied to sentiment analysis systems (Lakkaraju et al., 2024b) and then extended to composite NLP tasks (Lakkaraju et al., 2023) and time-series forecasting models (Lakkaraju et al., 2024a, 2025). These methods go beyond correlation by estimating treatment effects and adjusting for confounding. In this paper, we build on that foundation but shift the role of ratings: rather than treating them solely as evaluation metrics, we use them to explain model behavior. By integrating ratings with traditional XAI methods, we allow stakeholders to compare models, test hypotheses, and interpret robustness and bias in a more structured way.
2.4. Understanding Stakeholder Needs in XAI
Understanding stakeholder needs is critical for building explanations that are useful, not just technically accurate (Bhatt et al., 2020). For example, (Bundas et al., 2023) highlights NIST’s emphasis on hierarchical vocabularies: high-level concepts should be broadly accessible, while lower levels serve expert users. Studies show stakeholders prefer interactive explanations. Interviews in (Hoffman et al., 2023b) show a strong preference for tools that allow input manipulation and scenario exploration. RDE supports this by allowing users to perturb inputs and observe causal effects on model behavior (Figure 12). While current XAI methods focus on local explanations, users often want a broader understanding of model behavior (Mueller et al., 2019). (Suresh et al., 2021) proposes a three-level framework for interpretability needs, including the need to understand model limitations. RDE addresses this by showing how behavior changes under different conditions, supporting users with a structured robustness evaluation. In this paper, we show how H-XAI brings these ideas together, supporting interactive, stakeholder-aligned explanations through a combination of rating and post-hoc XAI methods.
3. Approach
Figure 2. RDE Workflow.
To explain model behavior in ways that support robustness, fairness, and practical decision-making, we propose a hybrid framework called Holistic-XAI (H-XAI). It integrates traditional post-hoc explanation methods with rating-driven explanation (RDE) approach. Unlike approaches that rely solely on post-hoc explanations, H-XAI combines both: traditional XAI methods are well-suited for explaining individual predictions and feature attributions, while RDE supports broader questions about robustness, bias, and perturbation effects. This combination lets stakeholders examine both how individual predictions are made and how the model behaves under different input conditions.
3.1. Rating-Driven Explanation (RDE)
RDE combines three components: (a) a causal model that specifies the structural assumptions between key attributes; (b) a set of metrics to quantify causal and statistical bias; and (c) a workflow that connects user queries to ratings and model comparisons.
3.1.1. Causal Model
We begin by specifying a causal model for the task, which includes assumptions about how attributes influence one another. Figure 1(a) shows the generalized causal diagram used across our case studies. While the causal model captures the underlying assumptions, the diagram is its visual representation, making explicit the directional relationships we test. In this setup, the outcome $O$ (e.g., prediction or residual error) is influenced by a treatment variable $T$ (e.g., credit amount, input perturbation), which may itself be influenced by a protected attribute $Z$ (e.g., gender and age). If $Z$ is a common cause of both $T$ and $O$, then confounding exists and must be accounted for to estimate a valid treatment effect.
This diagram allows us to define target causal estimands such as $E[O\mid do(T)]$ (i.e., the deconfounded distribution after adjustment), which represents the expected model outcome under an intervention on $T$. In practice, causal models are often specified by experts, derived from controlled experiments, or inferred using causal discovery algorithms (Gong et al., 2024; Varando et al., 2024), which themselves rely on statistical associations in the observed data. We begin with a hypothesized causal model that assumes protected attributes act as confounders and explore different scenarios within this structure. Even when based on assumptions, such models provide a structured framework for reasoning under interventions, something that correlational methods often fail to do, especially in the presence of confounding (as illustrated by Simpson’s paradox (Pearl, 2022)). For instance, in Section 4.1, we consider the possibility that a loan applicant’s status influences the credit amount requested, an entirely plausible scenario in practice. The outcome $O$ may be either the model prediction $\hat{Y}$ or a derived quantity g(Y^)g($\hat${Y}). For example, in time-series settings, we define $O=|\hat{Y}-Y|$ as the residual error. This lets us ask: “Does a certain input perturbation systematically worsen performance for a particular group?”.
3.1.2. Rating metrics
While RDE builds on prior rating methods, it repositions their function. Earlier works primarily used ratings to compare model performance across input conditions. RDE instead uses them as part of a structured explanation workflow: formulating a question (e.g., “Does treatment T affect outcome O?”), selecting a baseline, applying a metric, and interpreting the result as a potential explanation. This gives ratings an explanatory role, especially in helping stakeholders reason about model bias, instability, or robustness under different conditions. We use three metrics, each designed to answer a specific kind of question. The full implementation details are in Appendix A.
Weighted Rejection Score (WRS). Let $Z$ have $m$ groups and let $\mathcal{C}=\{c_{1},\dots,c_{K}\}$ be a set of confidence levels (e.g., 95%,75%,60%), with corresponding weights $w_{k}$. For each unordered pair $(a,b)$ of groups and each confidence level $c_{k}$ compute the two-sample $t$-test statistic $t_{ab}^{(k)}$ and define the indicator
$I_{ab}^{(k)}=\begin{cases}1&\text{if }|t_{ab}^{(k)}|>t_{\mathrm{crit}}(c_{k},\,\nu_{ab}^{(k)})\\ 0&\text{otherwise,}\end{cases}$
where $t_{\mathrm{crit}}(c_{k},\nu)$ is the critical $t$-value at confidence $c_{k}$ with degrees of freedom $\nu_{ab}^{(k)}$. The WRS is the weighted sum of these rejections:
$\mathrm{WRS}\;=\;\sum_{k=1}^{K}w_{k}\sum_{1\leq a<b\leq m}I_{ab}^{(k)}.$
In our experiments we use $\mathcal{C}=\{0.95,0.75,0.60\}$ with weights $w=\{1.0,0.8,0.6\}$; the inner sum runs over all $\binom{m}{2}$ group pairs.
Average Treatment Effect (ATE). When the question is about the effect of a specific change (e.g., halving the requested loan amount), we estimate the ATE,
$\text{ATE}\;=\;\mathbb{E}[O\mid\text{do}(T=1)]\;-\;\mathbb{E}[O\mid\text{do}(T=0)],$
which is the average change in the model outcome under the intervention. In practice we estimate this with propensity-score matching (PSM) or G-computation when the treatment is continuous. Small ATE means the model is effectively insensitive to that change; a large ATE means the change shifts predictions or errors.
Deconfounded Impact Estimation (DIE%). An observed ATE can be misleading if a protected attribute $Z$ affects both $T$ and $O$. To quantify that distortion we compute
$\text{DIE\%}\;=\;100\times\frac{\lvert\text{ATE}_{\text{unadj}}-\text{ATE}_{\text{adj}}\rvert}{\lvert\text{ATE}_{\text{adj}}\rvert}$,
where $\text{ATE}_{\text{adj}}$ is the ATE after adjusting for $Z$ (via PSM or G-computation) and $\text{ATE}_{\text{unadj}}$ is the ATE before adjusting. If DIE% is large, much of the apparent effect of $T$ is actually explained by group differences; if DIE% is small, the treatment effect is robust to confounding.
These three measures play distinct roles in explanation: WRS measures distributional disparities across groups; ATE quantifies how much an intervention changes model outputs on average; DIE% quantifies the effect of confounders on the relationship between treatment and the outcome. We selected these metrics because they are straightforward to explain to non-technical stakeholders (regulators, applicants) and because together they let us rapidly triage whether a model’s behavior warrants a deeper, instance-level analysis (for example, using SHAP or counterfactuals).
3.1.3. RDE Workflow
Figure 2 shows the end-to-end RDE process. Given a user query (e.g., “Is model performance consistent across demographic groups?”), we: 1. Define a treatment ($T$), outcome ($O$), and protected attribute ($Z$), 2. Select the appropriate metric (e.g., ATE or DIE %), 3. Compute scores for the target / test model, 4. Compare it to two automatically generated reference models: - A random model, which outputs predictions at random, - A biased model, which gives different, often extreme predictions based solely on the value of the protected attribute $Z$. For example, in the context of credit risk assessment, always classifying one group as good risk and another group as bad risk.
This comparison grounds the explanation. A model that behaves closer to the biased baseline may raise fairness concerns; one that matches the random model may be unreliable. These comparisons are not just for evaluation: they help users understand why the model’s behavior is problematic and where. RDE outputs can be numeric scores or visual plots, depending on the user query. For example, Figure 8 shows a case where RDE highlights age-based confounding in a credit classifier, while Figure 12 shows how error rates change under missing input perturbations.
3.2. Traditional XAI Methods Used
While RDE provides explanations concerning the robustness of the models at a global level, it is not suitable for local feature attribution or communicating the robustness of the model at a global level. To address these needs, we integrate standard post-hoc explanation methods, which stakeholders may already be familiar with.
Partial Dependency Plots (PDPs): Show the marginal effect of a feature on predictions by averaging over the dataset. Useful for understanding monotonicity or directionality, but unreliable when features are correlated. In Figure 8, we use PDPs to examine the effect of credit amount across applicant groups.
SHapley Additive exPlanations (SHAP): Estimates the marginal contribution of each feature to a prediction, using additive Shapley values. For time-series models, we use TreeSHAP on surrogate models (Raykar et al., 2023) to reduce computational cost. This makes the explanations more scalable, even for long input sequences. In our case study, we evaluated the surrogate on a small subset (10% of data, 60 samples) and found that its MASE and SMAPE differed from MOMENT’s by only 0.66 and 0.18%, respectively (see Table A.1 in the appendix).
Counterfactuals: Identify minimal input changes needed to flip a prediction (Wachter et al., 2017). This is useful for questions like, “What should this applicant change to get approved?”
3.3. Holistic eXplainable AI (H-XAI)
H-XAI brings together traditional post-hoc methods, some of which explain individual predictions (like SHAP and counterfactuals) and others that show general patterns (like PDPs or global SHAP), with RDE, which focuses on robustness by testing how model behavior changes under interventions and comparing it to biased and random baselines. This lets us analyze not just how the model works, but how stable and fair it is under different conditions. This process is often driven by stakeholders themselves, who have hypotheses involving certain variables they want to test. Stakeholders may begin with a general question (“Is the model fair across groups?”), use RDE to detect problems, then drill down into specific examples using SHAP or counterfactuals to understand the source of those issues. As shown in Figure 1(c), this setup allows users to move between different types of explanations depending on the question, without being locked into a single method or interpretability style. We demonstrate this combined framework through two use cases in Section 4, mapping specific stakeholder questions to the methods used to answer them.
Figure 3. (a) Causal diagram for binary classification on the German Credit dataset. Protected attributes include Age, Personal Status, Gender, or their combinations, depending on query. (b) Causal diagram for time-series forecasting (stock prices), adapted from (Lakkaraju et al., 2024a).
4. Case Studies
Figure 4. Hypothesis: Does decreasing the credit amount by 50% improve the chances of loan approval? Conclusion: On average, decreasing the requested credit amount by 50% increases the risk by 0.06, meaning applicants will be more likely to be classified as good risk (denoted by class 1). This suggests that asking for a smaller loan may improve the likelihood of loan approval only marginally.
Figure 5. Hypothesis: Does doubling applicants’ age, reduce their chances of being approved for a loan? Conclusion: On average, doubling a person’s age led to a noticeable drop in the predicted creditworthiness (risk score decreased by 0.35). This suggests that the model tends to classify older applicants as higher risk, which could be a sign of age-related bias.
Figure 6. Hypothesis: Does decreasing the loan duration by half improve the chances of loan approval? Conclusion: Reducing the loan duration by half led to a slight increase in predicted creditworthiness (risk score increased by 0.097). This indicates that the model tends to view shorter loan durations more favorably, which may slightly help an applicant get approved.
Figure 7. Stakeholder: Jack (applicant; individual); Scenario-1: Bank XYZ uses a Random Forest model to classify applicants as good or bad risk based on features like credit amount, age, personal status, gender, and more. Jack was classified as a bad risk and wants to understand why, and what changes he could make to be considered a good risk and get his loan approved.
Figure 8. Stakeholder: Maya (Regulatory official; Regulatory body). Scenario-2: An applicant reports that the AI model used by Bank XYZ is biased with respect to personal status and gender, and suspects their loan was rejected for this reason. Maya, a regulator from the ABC Bureau, begins an investigation to determine whether this bias is present.
Figure 9. Stakeholder: James (Data Scientist; Operational Organization). Scenario-3: Given a set of candidate models, James sets out to evaluate which one performs best on the client’s credit risk data. He plans to assess both accuracy and fairness before selecting a model for deployment; Interestingly, the global SHAP summary (Q2) is closely aligned with the Jack’s SHAP values shown in Figure 7!
Question Method Used Limitation When to Combine
Do outcomes differ across groups (e.g., gender, age)? WRS Cannot isolate cause of disparity (only statistical difference). Combine with DIE % to check if protected attribute has a confounding or direct effect.
Does a certain input (e.g., credit amount) causally influence the outcome? ATE Does not explain at the individual level; ignores confounding if unadjusted Use DIE% to validate causal effect is not confounded
Is the effect of a feature (e.g., credit amount) on the outcome distorted because it is influenced by a protected attribute (or confounding) (e.g., age)? DIE% Cannot be used when treatment is uniform; needs good causal estimation models; not instance-level Use SHAP to identify where and how confounding influences predictions at the instance level
How does a feature influence predictions overall? PDP Fails when features are correlated; only shows average trends Use ATE for more reliable estimation if confounding is present.
Which features contribute most to a single prediction? SHAP (local) Can be unstable with feature interactions; does not show global behavior Use global SHAP or ATE to see if local pattern is consistent across dataset
Which features matter globally? SHAP (global) Attribution unreliable under correlated inputs; may mis-calculate importance Use RDE to test if feature importance remains stable across subgroups or under input perturbations.
What minimal change would flip a decision? Counterfactual Explanation May suggest unrealistic or invalid inputs; ignores global trends Validate counterfactual’s plausibility using SHAP or PDP (helps identify features that matter to the model).
Does the model behave differently for one group under noisy/missing input? DIE% Needs a causal model. Use SHAP to identify which feature contributions shift across groups post-perturbation.
Is the model sensitive to a certain input change? ATE (perturbation as treatment) Only provides average effect; hides heterogeneity. Use SHAP to see variability of effect across instances.
Does this error pattern resemble biased or random behavior? RDE with biased/random baselines Baselines are synthetic; may not match real misuse cases. Combine with counterfactuals or SHAP to examine edge cases or deviations.
Table 1. Mapping explanation methods to stakeholder questions, limitations, and when to combine them. Combinations show scenarios from our case studies (Section 4). Others are possible depending on task and data context.
In this section, we present two case studies to demonstrate how H-XAI can address the explanation needs of different stakeholders (introduced in Section 2.1). These scenarios are structured around a taxonomy of questions that align with the XAI Question Bank (Liao et al., 2021). The mapping of our scenario questions to the categories of the XAI Question Bank is provided in the appendix (Table A.2). We show six scenarios: three for each stakeholder type in a given task.
4.1. Case Study-1: Binary Classification
In the first case study, we choose binary credit risk classification as it is a well-studied, critical domain where fairness, explanation, and model behavior have real-world consequences (Demajo et al., 2020; Kłosok et al., 2020; Melsom et al., 2022).
A. Data. German Credit Dataset (Frank, 2010) has been widely used in prior work on algorithmic fairness and explainability (Zemel et al., 2013; Martins et al., 2023; Verma and Rubin, 2018), making it suitable for examining how different stakeholder questions surface in a binary classification task. It contains 1,000 rows with 20 applicant features, including demographic information (e.g., age, personal status and sex), financial details (e.g., credit amount, duration, checking account status), and other applicant-related attributes which are used to assess loan risk. We selectively analyze subsets of these features depending on the specific explanation or causal question being addressed. The primary outcome variable is binary: $Risk\in\{0,1\}$, where $Risk=1$ indicates a good credit risk (leading to loan approval), and $Risk=0$ indicates a bad credit risk (loan rejection).
B. Causal Setup. For generating RDE, we use the causal diagram shown in Figure 3(a). Specifically, we focus on a smaller set of variables: we treat Credit Amount as the treatment, Age and Personal Status and Sex as protected attributes following (Zemel et al., 2013), and Risk as the outcome. In contrast, when applying XAI methods like SHAP, we do use the full feature set, since those methods are not bound by the same identifiability constraints and are designed to give feature attribution-based explanation rather than estimate causal effects.
C. AI Models Considered. We use two models: a linear, interpretable model (Logistic Regression), and a non-linear ensemble model (Random Forest). They are two of the most popular methods used for credit scoring (Addy et al., 2024). Figure 9 highlights an accuracy versus confounding bias trade-off in Scenario-3. While Random Forest outperforms Logistic Regression in accuracy, it also exhibits higher confounding bias. This shows the common, but not universal, trade-off between accuracy and interpretability (Atrey et al., 2025).
We present three scenarios in which H-XAI (described in Section 3) is applied to address questions posed by different types of stakeholders: an individual applicant (Figure 7), a regulatory official (Figure 8), and a data scientist within an operational organization (Figure 9).
We also provide an example of a hypothesis-driven explanation for credit risk classification in Figures 4, 5, 6. We define hypothesis-driven explanations as a structured process in which users test specific assumptions or causal questions about model behavior, rather than passively receiving fixed outputs. In our formulation, such hypotheses are operationalized through causal models, allowing users to quantify effects and systematically examine how the model behaves. While this is not exactly what Miller envisioned, it aligns with the direction of his Evaluative AI framework (Miller, 2023). The causal setup described for RDE in Section 3.1 also supports hypothesis-driven explanations, allowing users to test claims about model behavior rather than receive fixed outputs, as illustrated in Figure 4. We show how doubling age (Figure 5) or halving loan duration (Figure 6) affects predicted credit risk, demonstrating how hypothesis-driven explanations can serve as a standalone method for validating assumptions about model behavior.
Figure 10. Stakeholder: Suresh (Regulatory official; Regulatory body). Scenario-4: Suresh uses a stock forecasting tool developed by a third-party company to guide her investment decisions. He notices that the tool performs poorly on certain days and suspects it may be less reliable for some companies. To decide whether to continue using it, he wants to understand how robust the tool is.
Figure 11. Stakeholder: Jane (Retail Investor; Individual). Scenario-5: Jane uses a stock forecasting tool developed by a third-party company to guide her investment decisions. She notices that the tool performs poorly on certain days and suspects it may be less reliable for some companies. To decide whether to continue using it, she wants to understand how robust the tool is.
Figure 12. Stakeholder: Amanda (Data scientist; Operational organization). Scenario-6: Amanda is evaluating several time-series forecasting models for deployment. She wants to compare their accuracy and bias to choose one that best fits her team’s priorities.
4.2. Case Study-2: Time-Series Forecasting
In the second case study, we focus on stock price forecasting, a task that is high-impact in financial decision-making and widely used in algorithmic trading, portfolio management, and investment analytics.
A. Data. We collected daily stock prices from Yahoo! Finance for six companies across different industries: Meta (META) and Google (GOO) in social technology, Pfizer (PFE) and Merck (MRK) in pharmaceuticals, and Wells Fargo (WFC) and Citigroup (C) in financial services, as closing stock prices is the most commonly used data source for time-series forecasting (Lin and Marques, 2024; Lakkaraju et al., 2024a). The dataset covers from March 28, 2023, to April 22, 2024; data up to March 22, 2024, was used to forecast the following month.
B. Causal Setup. To generate RDE for time-series forecasting, we adopt the preprocessing approach from (Lakkaraju et al., 2024a), where the difference between the 20 predicted stock prices (based on the previous 80 time steps) and their corresponding ground truth values is used to compute a sequence of residuals. While other metrics like SMAPE or MASE are available, we use residuals because they give the raw prediction error at each point, rather than at an aggregate level. To capture worst-case performance, we take the maximum residual within each prediction window as the outcome of interest. For simplicity, we refer to this as the residual throughout the paper. While predicted stock prices may vary across companies, our focus is on whether residuals, i.e., model errors, also vary systematically, which may indicate unreliability. This residual serves as the basis for analyzing how protected attributes or perturbations affect model behavior. Figure 3(b) shows the causal diagram used for this analysis. We treat Perturbation (e.g., introducing missing values or zero-imputation, as in Figure 12) as the treatment, Company as the sensitive attribute, and Residual as the outcome. This causal setup is used exclusively for explanation questions that require RDE.
C. AI Models Considered. Our goal is not to benchmark forecasting performance but to analyze and explain model behavior. So rather than selecting the most competitive models, we chose two representative ones from the models used in (Lakkaraju et al., 2024a): MOMENT, an open-source, time-series-specific foundation model, and Gemini, a general-purpose foundation model, both of which are easily accessible. We also use ARIMA, a statistical time-series forecasting model, as an additional baseline.
We present three scenarios in H-XAI and apply them to address questions posed by different types of stakeholders: a regulatory official (Figure 10), a retail investor (Figure 11), and a data scientist within an operational organization (Figure 12).
4.3. When to Use: Mapping Questions to Methods
A key goal of H-XAI is to help users in choosing the right tool for the right question, not only to apply methods indiscriminately. Table 1 maps stakeholder questions to fitting methods. We summarize what each reveals, their limitations, and when they require complementary methods. This is where H-XAI matters: it lets users pose questions, pick methods, identify problems, and cross-validate them using complementary methods. No single method is sufficient; the strength lies in the combination.
5. Discussion and Conclusion
Although our case studies focus on credit risk classification and stock price forecasting, H-XAI is task-agnostic and extends beyond these domains. The framework centers on selecting explanation techniques based on stakeholder roles and the questions they bring, which is increasingly important as AI systems are deployed through web-based decision services and online platforms. In classification settings, RDE can estimate the causal impact of applicant attributes on model decisions; in forecasting settings, it can quantify how input perturbations affect a model’s stability. Traditional XAI methods complement RDE by providing instance-level explanations where causal analysis may not apply.
As shown in Figure 2, RDE is more than a collection of metrics: it provides a step-by-step reasoning process that involves constructing appropriate baselines, outlining a causal model, and choosing the corresponding estimands. This structure allows H-XAI to adapt to a variety of stakeholder needs, whether the goal is to investigate robustness, quantify bias, or interpret individual decisions. These use cases naturally arise in settings where users interact with AI predictions through online interfaces, financial dashboards, or automated decision portals. We recognize that stakeholders may not always know how to specify a causal model. To handle this, RDE adopts a generalized causal model by default, capturing structural relationships that appear across many tasks. Depending on how a question is framed, the analysis may target confounding bias (e.g., DIE%) or statistical bias (e.g., WRS), as illustrated in Q1 of Figure 7 versus Q1 of Figure 12. Throughout the framework, interpretability is not treated as an inherent property of a single method but as something that emerges from aligning the method with the stakeholder’s specific question and context.
A core insight from this work is that explanation is an iterative process rather than a one-off output. Different questions require different tools, and no individual method is adequate on its own. By combining RDE with counterfactual reasoning and feature-based explanations, we shift the role of explanation from post hoc justification to a form of structured interaction that allows stakeholders to test hypotheses and improve their understanding of model behavior. While this paper presents a first step toward a holistic approach, several directions remain open for future work, including additional case studies, exploring complementary explanation methods, and conducting user studies with diverse stakeholders who rely on AI-driven decisions in online and operational settings.
Appendix
Appendix A RDE Metrics
Here is how the rating metrics are calculated.
  • Algorithm 1 computes the Weighted Rejection Score (WRS) to measure statistical bias by analyzing how sensitive the model’s outcomes are to variations in sensitive attributes.
  • Algorithm 2 calculates Deconfounding Impact Estimation (DIE %) using PSM to assess confounding bias by measuring the effect of the confounder on model outcomes before and after deconfounding.
  • Algorithm 3 calculates ATE by evaluating the difference in the model’s outcomes for treated data (perturbation applied) and control data (no perturbation) to evaluate the impact of the treatment on the outcome. The ATE is computed after applying PSM or G-computation to mitigate the confounding effects, hence the use of the do(.) operator in the ATE formulation.
Algorithm 1 WeightedRejectionScore
Purpose: Calculate the weighted sum of null-hypothesis rejections for dataset $d_{j}$ and model $s$ using confidence intervals $CI_{k}$ and weights $w_{k}$.
Input:
$d$: Dataset for a specific perturbation;
CICI: Confidence intervals (e.g., 95%, 70%, 60%);
$s$: AI model from the set of test models $S$;
$W$: Weights for different CIs (e.g., 1, 0.8, 0.6)
Output: $\psi$: Weighted rejection score
[…]
2:for each $(ci_{i},w_{i})$ in $(CI,W)$ do
3:  for each pair $(z_{a},z_{b})$ in classes of $Z$ do
4:    $(t,pval,dof)\leftarrow\text{T-Test}(z_{a},z_{b})$
5:    $t_{crit}\leftarrow\text{LookUp}(ci_{i},dof)$
6:    if $t>t_{crit}$ then
7:     $\psi\leftarrow\psi+w_{i}$
8:    end if
9:  end for
10:end for
11:return $\psi$
Algorithm 2 ComputeDIEScore
Purpose: Calculate the Deconfounding Impact Estimation (DIE) using Propensity Score Matching (PSM).
Input:
$s$: AI model,
$d$: Dataset corresponding to a perturbation,
$p$: A perturbation other than $p_{0}$ (control or no perturbation)
Output: $\psi$: DIE score
[…]
4:return $\psi$
Algorithm 3 ComputeATEScore
Purpose: Calculate the Average Treatment Effect (ATE).
Input:
$s$: AI model,
$d$: Dataset,
$p$: Treatment perturbation,
$p_{0}$: Control (unperturbed baseline)
Output: $\psi$: ATE score
[…]
2:return $\psi$
Model SMAPE (%) MASE
Surrogate 7.88 4.77
MOMENT 8.06 5.45
Table A.1. MASE and SMAPE scores on a 10% evaluation subset (60 samples) for the MOMENT model and its surrogate. The small differences indicate that the surrogate closely approximates MOMENT’s predictions for explanation purposes.
[…]
black
Scenario Question in Paper Mapped Question from XAI Question Bank
Case Study 1: Binary Classification (Credit Risk)
[…]
Individual (Fig. 7) Q2: How did each factor… contribute to my loan rejection? Why: What feature(s) of this instance determine the system’s prediction of it
Individual (Fig. 7) Q3: What is the smallest change I could make to my loan duration… to get my loan approved? How to be that: What is the minimum change required for this instance to get a different prediction?
Individual (Fig. 7) Q4: What is the smallest change I could make to my checking account balance to get my loan approved? How to be that: What is the minimum change required for this instance to get a different prediction?
[…]
Data Scientist (Fig. 9) Q1: How can I rate the models that are available in my company for accuracy? Performance: How accurate are the predictions?
[…]
Data Scientist (Fig. 9) Q3: On average, how does each feature influence the predicted credit risk in the logistic regression model? How: How does the system make predictions?
Case Study 2: Time-Series Forecasting
Retail Investor (Fig. 11) Q1: What in the recent price trend suggests that Google’s stock will drop tomorrow? Why: What features of this instance determine the system’s prediction of it?
Retail Investor (Fig. 11) Q2: What would the forecast be if the stock price at the most influential past time step were \$100 lower? What If: What would the system predict if this instance changes to…?
[…]
Regulatory (Fig. 10) Q1: A user reported that the forecasting tool… relies heavily on older stock prices. Does the model overall, depend more on long-term patterns than recent ones? How: How does it weigh different features?
[…]
Table A.2. Mapping of all scenario questions to the XAI Question Bank from (Liao et al., 2021). The highlighted questions in the table are specifically focused on model fairness and sensitivity to perturbations. While the question bank lacks explicit questions for these topics, we have mapped our questions to the closest available ones, highlighting a need for the original taxonomy to be extended to include these topics.
Figures 8,9 show additional scenarios from case studies illustrated in Section 4. Table A.2 shows how the questions from the paper’s case studies align with an existing XAI taxonomy (Liao et al., 2021). The table highlights specific questions on model fairness and sensitivity to perturbations, which lack explicit counterparts in the question bank.